is it best to solve 11x + 17y = 94 and x + y = 7 with substitution, elimination, or graphing
Answer by richwmiller(17219)

Elimination would involve multiplying whole equations by 11 or 17
Graphing is not accurate
So I would solve by substitution.
x = 25/6, y = 17/6
11x + 17y = 94
x + y = 7
x=7-y
11(7-y) + 17y = 94
77-11y+17y=94
6y=17
y=17/6
x + y = 7
x + 17/6 = 7
x=(42-17)/6
x=25/6